# STC-ViT: Spatio Temporal Continuous Vision Transformer for Weather Forecasting

### v2

@@ 5 Experiments and Results
Dataset. We trained STC-ViT on ERA5 dataset at resolution of $5.625^{\circ}$ (32 x 64 grid points) and $1.40625^{\circ}$ (128 x 256 grid points) provided by WeatherBench (Rasp et al., 2020). Details of the dataset are provided in Appendix A.
@@
Training Details. We consider weather forecasting as a continuous spatio temporal forecasting problem i.e a tensor of shape $N\times H\times W$ at time $t$ is fed to the pre-processing layer of the model where it passes through pre-derivation to STC-ViT and outputs a tensor of $N^{\prime}\times H^{\prime}\times W^{\prime}$ at future time step $t+\triangle t$. Complete training details of the model are given in the in Appendix B.
Normalization Experimental analysis show that z-score normalization produces superior results when applied to atmospheric variables such as geopotential at 500hpa pressure (z500) and temperature at 850hpa pressure (T850) and min-max produces better results when applied to surface variables such as 2-meter temperature (T2m) and mean sea-level pressure (MSLP). We compare the results of both methodologies at lead times of 1,3,5,7 and 10 days. The RMSE scores and comparisons are shown in figure 4.
Figure 4: The results comparison of normalized derivatives show that z-score normalization performs better for atmospheric variables while min-max produce better results for static variables
Ablation Studies We performed extensive empirical analysis to understand the role of different components in our model’s performance. The results of ablation studies in figure 5 show that while Neural ODE is a powerful architecture, it outperforms vanilla architectures when embedded in attention mechanism. Details of ablation studies are given in Appendix B.6.
Figure 5: Ablation studies highlight the results for different components of STC-ViT individually and show that the continuous attention component alone performs better than all other even NODE. Thus combined with NODE outperforms state-of-the-art results.
@@
Results and Discussion We trained STC-ViT on ERA5 dataset at resolutions of $5.625^{\circ}$ for lead times ranging from 6 to 36 hours and $1.40625^{\circ}$ for lead times ranging from 1 to 10 days and call them $STC_{5.6}$ and $STC_{1.4}$ respectively. We compare $STC_{5.6}$ forecast results with ClimaX, FourCastNet and ClimODE on ERA5 dataset at $5.625^{\circ}$ resolution shown in Table 1. $STC_{5.6}$ outperforms ClimaX, FourCastNet and ClimODE at all lead times which shows that replacing regular attention with continuous attention in ViT architecture derives improved feature extraction by mapping the changes occurring between successive time steps.
Additionally, we compare the performance of $STC_{1.4}$ with the results of GraphCast and PanguWeather which are trained at a higher resolution of $0.25^{\circ}$ shown in Figure 7. $STC_{1.4}$ outperforms PanguWeather for medium-range forecasting starting at 3-day lead time and is on par with GraphCast. The superior performance of GraphCast and STC-ViT also opens a venue for future research in which image data could be modeled as graph and then processed through continuous attention mechanism. We also compare our results with both versions of IFS Andersson (2022), IFS-HRES which is state-of-the-art forecasting model at high resolution of $0.1^{\circ}$ and IFS-ENS, ensemble version trained at $0.2^{\circ}$.
Figure 6: RMSE comparison of STC-ViT trained at $1.40625^{\circ}$ with GraphCast and PanguWeather trained at $0.25^{\circ}$, IFS-ENS at $0.2^{\circ}$ and IFS-HRES at $0.1^{\circ}$ resolution data for lead times ranging from 1 to 10 days
Figure 7: RMSE results of $STC_{5.6}$ compared against ClimODE (ODE based), ClimaX (non-pretrained) and FourCastNet (FCN) trained on ERA5 at $5.625^{\circ}$ resolution
@@
Table 1: RMSE and ACC results of $STC_{5.6}$ compared against ClimODE (ODE based), ClimaX (non-pretrained) and FourCastNet (FCN) trained on ERA5 at $5.625^{\circ}$ resolution
@@
RMSE (Lower is better) ACC (Higher is better) Variable Lead Time (hrs.) STC-ViT ClimODE ClimaX FCN STC-ViT ClimODE ClimaX FCN z500 $(m^{2}\backslash s^{2})$ 6 59.61 102.9 247.5 149.4 1.0 0.99 0.97 0.99 12 76.20 134.8 265.3 217.8 1.0 0.99 0.96 0.99 18 86.10 162.7 319.8 275.0 1.0 0.98 0.95 0.99 24 92.09 193.4 364.9 333.0 1.0 0.98 0.93 0.99 36 111.31 259.6 455.0 449.0 0.99 0.96 0.89 0.99 T2m (K) 6 0.89 1.21 2.02 1.28 1.00 0.97 0.92 0.99 12 0.92 1.45 2.26 1.48 0.99 0.96 0.90 0.99 18 0.96 1.43 2.45 1.61 0.99 0.96 0.88 0.99 24 1.04 1.40 2.37 1.68 0.98 0.96 0.89 0.99 36 1.09 1.70 2.87 1.90 0.98 0.94 0.83 0.99 T850 (K) 6 0.84 1.16 1.64 1.18 1.00 0.97 0.94 0.99 12 0.88 1.32 1.77 1.47 1.00 0.96 0.93 0.99 18 0.92 1.47 1.93 1.65 0.99 0.96 0.92 0.99 24 1.11 1.55 2.17 1.83 0.99 0.95 0.90 0.99 36 1.19 1.75 2.49 2.21 0.98 0.94 0.86 0.99 U10 $(m\backslash s)$ 6 0.98 1.41 1.58 1.47 1.00 0.91 0.92 0.95 12 1.06 1.81 1.96 1.89 0.99 0.89 0.88 0.93 18 1.17 1.97 2.24 2.05 0.99 0.88 0.84 0.91 24 1.29 2.01 2.49 2.33 0.98 0.86 0.80 0.89 36 1.34 2.25 2.98 2.87 0.98 0.83 0.69 0.85
@@ 6 Conclusion and Future Work
In this paper, we present STC-ViT, a novel technique designed to capture continuous dynamics of weather data. STC-ViT achieves competitive results in weather forecasting which shows that vision transformers can model continuous nature of spatio-temporal dynamic systems with carefully designed attention mechanism.
While STC-ViT outperforms various state-of-the-art weather forecasting models, it is important to address its limitations in weather forecasting systems. STC-ViT is a deterministic model which does not account for model uncertainty which can produce unrealistic results. Extending STC-ViT to a probabilistic model can be addressed in future works. Also, scaling STC-ViT to better accommodate the multi-modal high resolution training and evaluation presents an opportunity as future research. Finally, addressing the black-box problem of STC-ViT can shed light on model learning insights which is equally important for climate science community.
@@
Acknowledgements
This work was supported by SmartSAT Cooperative Research Center (project number P3-31s). We would also like to acknowledge NCI (National Computational Infrastructure) for providing us with GPU resources which enabled us to perform this research.
Appendix A Data
We used WeatherBench Rasp et al. (2020) to download the ERA5 Hersbach et al. (2018, 2020) dataset provided by the European Center for Medium-Range Weather Forecasting (ECMWF) for training our model. ERA5 is a weather reanalysis dataset of hourly intervals that provides data over several parameters such as land, ocean and atmosphere. The real-time data with the highest resolution of $0.25^{\circ}\times 0.25^{\circ}$ is being collected since 1940 till date. ERA5 is considered to be the most physically accurate dataset as it is based on data assimilation methodology in which new forecasts are continuously combined with previous forecasts in such a way that resultant is consistent with laws of physics. WeatherBench provides regridded data at resolutions of $5.625^{\circ}$ , $2.8125^{\circ}$ and $1.40625^{\circ}$. Table 2 provides a detail of weather variables used for training STC-ViT. For our experiments we split the data into training (1979-2017), validation (2018) and test (2019-2021) sets.
Table 2: ERA5 dataset variables and their atmospheric pressure values
Variable name (Type) Abbreviation Levels Geopotential (Atmospheric) Z 50 250 500 600 700 850 925 U Wind Component (Atmospheric) U 50 250 500 600 700 850 925 V Wind Component (Atmospheric) V 50 250 500 600 700 850 925 Temperature (Atmospheric) T 50 250 500 600 700 850 925 Specific Humidity (Atmospheric) Q 50 250 500 600 700 850 925 Relative Humidity (Atmospheric) R 50 250 500 600 700 850 925 Land-sea mask (static) LSM Orography (static) 2 meter Temperature (static) T2m Mean Surface level Pressure (static) MSLP 10 meter U Wind Component (static) U10 10 meter V Wind Component (static) V10
Appendix B Experiment Details
B.1 Evaluation Metrics.
@@ Appendix B Experiment Details
B.2 Optimization
We use the AdamW optimizer Loshchilov and Hutter (2017) with a learning rate of 0.001. We utilize a 10% linear warm-up, after which a Cosine learning rate scheduler is adopted which progressively lowers the learning rate to zero following the warm-up period. We train STC-ViT for 50 epochs which takes about 4-5 days in total.
B.3 Hyperparameters
@@
N Number of variables 48
p Patch size 4
@@
Depth Number of Transformer layers 8
@@
batch_size Batch Size 4
@@
ES rate Early stopping tolerance 20
B.4 Normalization
@@
B.5 Software and Hardware Requirements
We use PyTorch Paszke et al. (2019), Pytorch Lightning Falcon (2019), torchdiffeq Chen et al. (2018), and xarray Hoyer and Hamman (2017) to implement our model. We use 8 DGXA100 devices with 80GB VRAM for training STC-ViT. All architectural and training details are provided at https://anonymous.4open.science/r/STC-ViT-3276.
B.6 Ablation Studies
Continuous Attention only model. To understand the importance of continuous attention, we remove the neural ODE layer and provide the output of attention mechanism to the basic feed forward network. While continuous attention only model does not outperform the STC-ViT, it provides better prediction accuracy than the traditional attention model.
Continuous Attention vs Neural ODE layer. We perform another ablation study to compare how well Neural ODE capture continuity in weather information when simply added as a layer in the transformer architecture. Neural ODE Chen et al. (2018) is a continuous depth neural network designed to learn continuous information in the process as well. For this study, we replace the continuous attention with vanilla attention and add a Neural ODE layer in the feed forward block. The feed forward solution is approximated using Runge-Kutta (RK) numerical solver.
Vanilla Vision Transformer Vision Transformer has emerged as a powerful architecture which captures long-term dependencies better than any model. However, this ablation study shows the limitation of ViT in representing the dynamical systems such as weather forecasting. For this ablation study, we simply train a basic ViT architecture on ERA5 dataset. Compared with STC-ViT, ViT under performs for prediction accuracy showing the superiority of continuous models in weather forecasting systems.
Appendix C Results Summary
Figure 8: 1 day forecasts
Figure 9: 3 day forecasts
Figure 10: 5 day forecasts

### v3

@@ 5 Experiments and Results
5.1 Dataset
We train STC-ViT on ERA5 dataset Hersbach et al. (2018; 2020) provided by the European Center for Medium-Range Weather Forecasting (ECMWF). We compare STC-ViT against several weather forecasting models by training it at two different resolutions of $5.625^{\circ}$ (32 x 64 grid points) provided by WeatherBench (Rasp et al., 2020) and $1.5^{\circ}$ (121 x 240 grid points) provided by WeatherBench2 (Rasp et al., 2024).
5.2 Training Details.
@@ 5.1 Dataset
We consider weather forecasting as a continuous spatio temporal forecasting problem i.e a tensor of shape $N\times H\times W$ at time $t$ is fed to the pre-processing layer of the model where it passes through pre-derivation to STC-ViT and outputs a tensor of $N^{\prime}\times H^{\prime}\times W^{\prime}$ at future time step $t+\triangle t$. Complete training details of the model are given in the in Appendix A.
5.3 WeatherBench
We train STC-ViT on hourly data with following set of variables: Land Sea Mask (LSM), Orography, 10-meter U and V wind components (U10 and V10) and 2-meter temperature (T2m) in addition to 6 atmospheric variables: geopotential (Z), temperature (T), U and V wind components, specific humidity (Q) and relative humidity (R) at 7 pressure levels: 50 250 500 600 700 850 925. We use data from 1979-2015 for training, 2016 for validation and 2017-2018 for testing phase. We compare STC-ViT with ClimaX, and ClimODE on ERA5 dataset at $5.625^{\circ}$ resolution provided by WeatherBench (Rasp et al., 2020). To ensure fairness, we retrained ClimaX from scratch without any pre-training.
@@ 5.3 WeatherBench
STC-Vit outperforms ClimaX and ClimODE at all lead times which shows that replacing regular attention with continuous attention in ViT architecture derives improved feature extraction by mapping the changes occurring between successive time steps. Additionally, enforcing physical constraints in the model lead to improved prediction scores. The RMSE and ACC results are shown in Table 1.
@@
Table 1: RMSE and ACC results of $STC_{5.6}$ compared against ClimODE (ODE based) and ClimaX (non-pretrained) trained on ERA5 at $5.625^{\circ}$ resolution
@@
RMSE (Lower is better) ACC (Higher is better) Variable Lead Time (hrs.) STC-ViT ClimODE ClimaX STC-ViT ClimODE ClimaX z500 $(m^{2}\backslash s^{2})$ 6 78.36 102.2 249 0.99 0.99 0.97 12 99.42 132.7 265.3 0.99 0.99 0.96 18 117.16 163 319.8 0.99 0.98 0.95 24 141.45 193.4 455 0.98 0.98 0.89 36 206.59 259.6 455 0.97 0.96 0.89 U10 $(m\backslash s)$ 6 0.92 1.44 1.58 0.97 0.91 0.92 12 1.11 1.80 1.96 0.96 0.89 0.88 18 1.28 1.97 2.24 0.95 0.88 0.84 24 1.46 2.00 2.49 0.93 0.87 0.80 36 1.89 2.25 2.95 0.89 0.83 0.70 V10 $(m\backslash s)$ 6 0.95 1.53 1.60 0.97 0.92 0.92 12 1.15 1.81 1.97 0.96 0.89 0.88 18 1.32 1.95 2.26 0.94 0.88 0.83 24 1.50 2.02 2.48 0.93 0.86 0.80 36 1.92 2.29 2.94 0.88 0.83 0.70 T2m (K) 6 0.87 1.20 2.02 0.98 0.97 0.92 12 1.04 1.44 2.26 0.97 0.96 0.90 18 1.13 1.42 2.45 0.97 0.96 0.88 24 1.18 1.40 2.37 0.97 0.96 0.89 36 1.42 1.70 2.85 0.96 0.94 0.84 T850 (K) 6 0.83 1.16 1.64 0.98 0.97 0.94 12 0.99 1.32 1.77 0.97 0.96 0.93 18 1.09 1.47 1.93 0.97 0.96 0.92 24 1.19 1.55 2.17 0.97 0.95 0.90 36 1.44 1.75 2.49 0.95 0.94 0.86 u50 $(m\backslash s)$ 6 1.42 2.01 2.05 0.98 0.92 0.92 12 1.75 2.56 2.68 0.98 0.91 0.90 18 1.90 2.89 2.92 0.96 0.90 0.89 24 2.06 2.97 3.01 0.96 0.90 0.89 36 2.37 3.03 3.12 0.94 0.88 0.87 v50 $(m\backslash s)$ 6 1.45 2.01 2.12 0.95 0.90 0.90 12 1.73 2.20 2.32 0.94 0.89 0.88 18 1.89 2.27 2.31 0.92 0.87 0.85 24 2.00 2.35 2.41 0.91 0.87 0.84 36 2.19 2.51 2.59 0.90 0.85 0.83
5.4 WeatherBench2
To keep training consistent with WeatherBench 2, we utilize the training data from 1979 to 2018, validation data from 2019, and test data from 2020 year. We train STC-ViT on 6 hourly data for following variables: T2m, u10 and v10 wind components and mean sea-level pressure (MSLP) along with five atmospheric variables: geopotential height (Z), temperature (T), U and V wind components, and specific humidity (Q). These atmospheric variables are considered at 13 pressure levels: 50, 100, 150, 200, 250, 300, 400, 500, 600, 700, 850, 925, 1000 hPa. We also compare our results with both versions of IFS Andersson (2022), IFS-HRES which is state-of-the-art forecasting model at high resolution of $0.1^{\circ}$ and IFS-ENS, ensemble version trained at $0.2^{\circ}$. Additionally, we compare STC-ViT against GraphCast and PanguWeather which are trained at a higher resolution of $0.25^{\circ}$ and finally with Neural GCM trained at $0.7^{\circ}$.
STC-ViT shows competitive performance, outperforming Pangu and GraphCast particularly for geopotential, temperature and wind variables. This is due to the physics loss penalizing the model for not obeying physics and encouraging outputs that balance temperature evolution and respect thermodynamic laws by penalizing deviations from these energy fluxes. We find out that Neural GCM is the best performing model on atmospheric variables which could be due to the physics embedded in the dynamical core. We believe that STC-ViT could also benefit from high resolution training and physics directly incorporated in the model architecture which we will explore in future. The RMSE and ACC results are shown in Figures 2 and 4 respectively.
Figure 2: RMSE comparison of STC-ViT trained at $1.5^{\circ}$ with GraphCast, PanguWeather, Neural GCM, IFS-ENS at $0.2^{\circ}$ and IFS-HRES at $0.1^{\circ}$ resolution data for lead times ranging from 1 to 10 days
Figure 3: ACC comparison of STC-ViT trained at $1.5^{\circ}$ with GraphCast, PanguWeather, Neural GCM, IFS-ENS at $0.2^{\circ}$ and IFS-HRES at $0.1^{\circ}$ resolution data for lead times ranging from 1 to 10 days
5.5 Ablation Studies
Vanilla Vision Transformer Vision Transformer has emerged as a powerful architecture which captures long-term dependencies better than any model. For this ablation study, we simply train a basic ViT architecture on ERA5 dataset. Compared with STC-ViT, ViT under performs for prediction accuracy showing the superiority of continuous models in weather forecasting systems.
Continuous Attention. To understand the importance of continuous attention, we remove the neural ODE layer and provide the output of attention mechanism to the basic feed forward network. While continuous attention only model does not outperform the STC-ViT, it provides better prediction accuracy than the traditional attention model.
Continuous Attention + Neural ODE layer. We perform another ablation study to compare how well Neural ODE capture continuity in weather information when simply added as a layer in the transformer architecture. Neural ODE Chen et al. (2018) is a continuous depth neural network designed to learn continuous information in the process as well. For this study, we replace the continuous attention with vanilla attention and add a Neural ODE layer in the feed forward block. The feed forward solution is approximated using Runge-Kutta (RK) numerical solver.
Figure 4: Ablation studies highlight the results for different components of STC-ViT individually and show that the continuous attention component alone performs better than all other even NODE. Thus combined with NODE outperforms state-of-the-art results
@@ 6 Conclusion and Future Work
In this paper, we present STC-ViT, a novel technique designed to capture continuous dynamics of weather data while obeying the fundamental physical laws of Earth’s atmosphere. STC-ViT achieves competitive results in weather forecasting which shows that vision transformers can model continuous nature of spatio-temporal dynamic systems with carefully designed attention mechanism.
While STC-ViT performs competitively with state-of-the-art data-driven weather forecasting models, it is important to address its limitations in weather forecasting systems. STC-ViT is a deterministic model which does not account for model uncertainty which can produce unrealistic results. Additionally, predictions on longer lead times result in blurry forecasts. Extending STC-ViT to a probabilistic model can be addressed in future works. Also, scaling STC-ViT to better accommodate the multi-modal high resolution training and evaluation presents an opportunity as future research. Finally, addressing the black-box problem of STC-ViT can shed light on model learning insights which is equally important for climate science community.
@@
Appendix A Experiment Details
A.1 Evaluation Metrics.
@@ Appendix A Experiment Details
A.2 Optimization
We use the AdamW optimizer Loshchilov & Hutter (2017) with a learning rate of 5e-5. We utilize a CosineAnnealing learning rate scheduler is adopted which progressively lowers the learning rate to zero following the warm-up period which is 10% of the total epochs. We train STC-ViT for 50 epochs for $5.625^{\circ}$ and 100 epochs for $1.5^{\circ}$ resolution data .
A.3 Hyperparameters
@@
p Patch size 2
@@
Depth Number of Transformer layers 4
@@
batch_size Batch Size 12 for $5.625^{\circ}$, 1 for $1.5^{\circ}$
@@
ES rate Early stopping tolerance 10
$\alpha$ Kinetic Loss weight factor 0.1-0.5
$\beta$ Potential Loss weight factor 0.1-0.5
$\gamma$ Thermodynamic Loss weight factor 0.8
A.4 Normalization
@@
A.5 Software and Hardware Requirements
We use PyTorch Paszke et al. (2019), Pytorch Lightning Falcon (2019), torchdiffeq Chen et al. (2018), and xarray Hoyer & Hamman (2017) to implement our model. We use 2 NVIDIA DGX A100 devices with 80GB RAM for training STC-ViT at $1.5^{\circ}$ and 4 NVIDIA Tesla Volta V100-SXM2-32GB for the training at resolution of $5.625^{\circ}$.